OpenStudy (aaronandyson):

One pipe can fill a cistern in 3 hours less than the other.The two pipes together can fill the cistern in 6 hours 40 minutes.Find the time each pipe will take to fill the cistern @sepeario

OpenStudy (sepeario):

assume the two pipes are x and y. Hence \[x=y+3\] and \[x+y=400\] (I have converted 6hrs 40 minutes into 400 minutes)

OpenStudy (sepeario):

Can you use simultaneous equations to solve this?

OpenStudy (aaronandyson):

????????

OpenStudy (sepeario):

From the second equation, we can also find that \[x=400-y\]

OpenStudy (sepeario):

by rearranging the values, yes?

OpenStudy (sepeario):

You can then substitute that value for x into the first equation, leaving only one variable in the equation. This is known as solving through simultaneous equations.

OpenStudy (aaronandyson):

No 6hrs 40 mins will be 6 + 40/60 As per this book :/

OpenStudy (sepeario):

Yeah but I like to simplify it into a mixed number.. It's just easier working with that.

OpenStudy (sepeario):

Anyway let me complete the equation for you. \[400-y=y+\] and then \[2y=400, y=200\] If y = 200 minutes, then converted back to hours, it is 3 hours and 20 minutes.

OpenStudy (aaronandyson):

It's suppose to be 13 hrs and 15 hrs :/ @perl @Sepeario

OpenStudy (aaronandyson):

12hrs and 15 hrs**

OpenStudy (aaronandyson):

1/x + 1/x-3 = 20/3?? @perl

OpenStudy (sepeario):

Never mind, I stuffed up :/

OpenStudy (sepeario):

Sorry lol

OpenStudy (aaronandyson):

u mean r^1 and r^2 right??

OpenStudy (perl):

they are the unknown rates of the pipes

OpenStudy (aaronandyson):

what r subscripts ??????

OpenStudy (perl):

its a way to label our unknown rates

OpenStudy (perl):

you can use x and y if you like instead

OpenStudy (aaronandyson):

1 = (2x + 3)(20/3)????

OpenStudy (aaronandyson):

what next explain please

OpenStudy (perl):

i think we should do this over again

OpenStudy (aaronandyson):

Alright @perl

OpenStudy (perl):

lets do this carefully

OpenStudy (aaronandyson):

Can't we do 1/x + 1/x-3 = 20/3?

OpenStudy (perl):

yes that works

OpenStudy (perl):

i used t, and you used x thats the only difference

OpenStudy (perl):

Let x and y be the rates of the first and second pipe respectively. One pipe can fill a cistern in 3 hours less than the other. I am going to use : distance = rate * time, but distance is number of cisterns completed (like number of miles completed by a runner) Lets assume the time it takes for the first pipe with rate x to fill one cistern is t. Then the time it takes for the second pipe with rate y is t - 3. 1 cistern = x * t 1 cistern = y * (t - 3) 1 = x*t 1 = y ( t - 3) x = 1/t y = 1/ ( t - 3) Now it says if we have the two pipes working together, they can complete one cistern in 6 and 40/60 hour 1 = ( x + y ) ( 60 + 40/60) substitute 1 = ( 1/t + 1/(t-3) ) ( 60 + 40/60) 1 = ( 1/t + 1/(t-3) ) ( 20/3) 3/20 = 1/t + 1 / (t-3)

OpenStudy (aaronandyson):

1/x + 1/x-3 = 20/3 x-3+x/x^2-3x = 20/3?

OpenStudy (aaronandyson):

20x^2 - 66x + 9 = 0?

OpenStudy (aaronandyson):

That's what I wrote in a comment above. so now?

OpenStudy (perl):

theres an arithmetic mistake

OpenStudy (perl):

\(\color{blue}{\text{Originally Posted by}}\) @AaronAndyson 1/x + 1/x-3 = 20/3 x-3+x/x^2-3x = 20/3? \(\color{blue}{\text{End of Quote}}\) this is wrong it should be 1/x + 1/(x-3) = 3/20

OpenStudy (aaronandyson):

?

OpenStudy (perl):

your equation is incorrect

OpenStudy (perl):

correct equation: 1/x + 1/(x-3) = 3/20

OpenStudy (perl):

1/x + 1/(x-3) = 3/20 multiply both sides by x*(x-3) [x*(x-3)] (1/x + 1/(x-3)] =x*(x-3) 3/20 distribute x*(x-3)*1/x + x(x-3)*1/(x-3) =x*(x-3) 3/20 cancel (x-3) + x = x(x-3) * 3/20 2x -3 = (x^2 -3x) * 3/20 Multiply both sides by 20 20(2x-3) = (x^2 -3x)*3 40x - 60 = 3x^2 - 9x 3x^2 - 9x - 40x + 60 = 0 3x^2 - 49x + 60 = 0 (3x -4) ( x-15) = 0 x = 15, 4/3

OpenStudy (aaronandyson):

"(3x -4) ( x-15) = 0 " How???????????

OpenStudy (perl):

I factored 3x^2 -49x + 60 = 0

OpenStudy (uri):

omg, im so good with these, but @perl is better. >.<

OpenStudy (aaronandyson):

Explain the factoring process @perl

mathslover (mathslover):

I will make that look better for you. I hope Perl will not mind it. :) \[\cfrac{1}{x} + \cfrac{1}{x-3} = \cfrac{3}{20} \\ \cfrac{1}{x} \times x(x-3) + \cfrac{1}{x-3} \times x(x-3) = \cfrac{3(x)(x-3)}{20} \\ x-3 + x = \cfrac{3(x)(x-3)}{20} \\ 2x - 3 = \cfrac{3x(x-3)}{20} \\ (2x -3)(20) = 3x(x-3) \\ 40 x - 60 = 3x^2 -9x \\ 3x^2 -49x + 60 = 0 \\ 3x^2 - 45x - 4x + 60 = 0 \\ 3x(x-15) - 4(x-15) = 0 \\ (3x-4) (x-15) = 0 \\ 3x = 4 \ \text{OR} \ x = 15 \\ x = 4/3 \ \text{OR} \ x = 15 \] Just thought to include LaTeX to make the equations look attractive ;)

OpenStudy (perl):

3x^2 -49x + 60 = 0 We can break down the middle number -49x , then factor by grouping 3x^2 - 4x -45x + 60 = 0 x( 3x - 4) - 15( 3x - 4) = 0 (3x-4) ( x - 15) = 0

OpenStudy (perl):

you can also use quadratic formula if you're stumped how to factor the quadratic, or it does not factor at all (that can happen)

mathslover (mathslover):

And he actually found those numbers by : product of roots should be 3 * 60 = 180 sum of roots should be -49 and so, by guessing the roots, he found the two numbers that satisfy these two conditions are : -45 and -4 :)

OpenStudy (aaronandyson):

Thanks. now I need help with 2 more questions like this one

OpenStudy (perl):

whats strange about this problem, there are two solutions?

OpenStudy (aaronandyson):

Yes.

OpenStudy (perl):

oh wait, one of the solutions do not work

OpenStudy (perl):

we found that x = 15, and x = 4/3 but if you plug in x = 4/3 you get a problem, (a negative rate for the second pipe). so the only solution that works is x = 15

OpenStudy (perl):

@mathslover \(\color{blue}{\text{Originally Posted by}}\) @mathslover And he actually found those numbers by : product of roots should be 3 * 60 = 180 sum of roots should be -49 and so, by guessing the roots, he found the two numbers that satisfy these two conditions are : -45 and -4 :) \(\color{blue}{\text{End of Quote}}\) i think theres an issue here. $$ \Large { ax^2 + bx + c = 0 \\~\\ \text{sum of roots } r_1, r_2 \\ r_1\cdot r_2 = \frac{c}{a} \\ r_1 + r_2 = \frac{-b}{a} } $$

OpenStudy (perl):

what im saying is, the product of the roots is not a*c

mathslover (mathslover):

Pardon me for that, perl! Didn't get it right. Thanks for pointing it out.

OpenStudy (perl):

actually this isn't bad, we can go back and derive an interesting result what if we were to do something

OpenStudy (perl):

\[ \Large { ax^2 + bx + c = 0 \\~\\ \text{sum of roots } r_1, r_2 \\ r_1\cdot r_2 = \frac{c}{a} \\ r_1 + r_2 = \frac{-b}{a} } \] multiply both sides of first equation by a^2, second equation by a \[ \Large { \\ a^2 \cdot r_1\cdot r_2 = a^2 \cdot \frac{c}{a} \\ a \cdot( r_1 + r_2) = a \cdot \frac{-b}{a} }\] now rearrange and simplify \[ \Large { \\ (a \cdot r_1)(a \cdot r_2) = ac \\ a \cdot r_1 + a \cdot r_2 = -b }\] Now we have the form \[ \Large { \\ u\cdot v = ac \\ u + v = -b }\] So we see that if we can find two numbers that multiply to the product of the first and last number, a*c and add up to the middle number -b, we can factor the original quadratic.

OpenStudy (perl):

that negative sign in front of the b is bugging me ...

mathslover (mathslover):

That looks neat. Why does that negative sign bug you? Does it create any issues with the mathematical point of view? :O